# Advancing Glitch Classification in Gravity Spy: Multi-view Fusion with Attention-based Machine Learning for Advanced LIGO’s Fourth Observing Run

### v1

@@ 5 Conclusion
In this study, we present a novel classifier for the Gravity Spy project, specifically designed to enhance the accuracy of glitch classification within the LIGO gravitational-wave detectors. The proposed classifier demonstrates exceptional performance compared to the previous classifiers. Several factors contribute to the classifier’s distinctiveness. Leveraging the inception residual block allows for deeper architecture exploration, facilitating the extraction of discriminative glitch features. The intermediate fusion strategy captures cross-view correlations more effectively. Label smoothing addresses noisy training labels, enhancing the generalization of glitch predictions. The attention mechanism not only enhances model performance but also provides interpretability, potentially aiding volunteers and experts in understanding glitch classification decisions.
The new classifier has successfully been integrated into the ongoing O4 run, actively employed by both volunteers on the Zooniverse platform and LIGO experts. Its capacity to potentially help volunteers identify and categorize new glitch classes fits well with the larger context of improved detector sensitivity, where a more advanced model becomes essential for understanding the details of more intricate glitch patterns.
Acknowledgements
The authors extend their gratitude to the community-science volunteers of the Gravity Spy project, whose dedicated efforts played a pivotal role in bringing this ambitious project to fruition. The authors also thank ManLeong Chan for comments that improved the quality of this manuscript. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. Gravity Spy is supported by NSF grants IIS-2106882, 2106896, 2107334, and 2106865. CPLB acknowledges support from Science and Technology Facilities Council (STFC) grant ST/V005634/1. ZD acknowledges support from the CIERA Board of Visitors Research Professorship and NSF grant PHY-2207945. This document has been assigned LIGO document number LIGO-P2300458.
@@ Acknowledgements
The datasets from the Gravity Spy project, including machine learning and volunteer classification information, can be accessed through the Zenodo repositories glanzer2023data ; zevin2022gravity . The corresponding codes are also accessible on GitHub through Gravity Spy.

### v2

@@ 5 Conclusion
In this study, we present a novel classifier for the Gravity Spy project specifically designed to enhance the accuracy of glitch classification for the LIGO gravitational-wave detectors. Several factors contribute to the classifier’s distinctiveness. Leveraging the inception residual block allows for deeper architecture exploration, facilitating the extraction of discriminative glitch features. The intermediate fusion strategy captures cross-view correlations more effectively. Label smoothing addresses noisy training labels, enhancing the generalization of glitch predictions. The attention mechanism not only enhances model performance but also provides interpretability, potentially aiding volunteers and experts in understanding glitch classification decisions. Our tests demonstrate that the proposed classifier has exceptional performance compared to the previous classifiers.
The new classifier has successfully been integrated into analysis of the ongoing O4 run, and is actively employed by both volunteers on the Zooniverse platform and LIGO experts. Its capacity to potentially help volunteers identify and categorize new glitch classes aligns with the Gravity Spy project’s larger aim of improving detector sensitivity. A more advanced model is essential for understanding the details of more intricate glitch patterns. Following investigations of new glitch types [24, 21], the new classifier will be retrained with expanded training data for use in future detector-characterization studies.
Acknowledgments
The authors extend their gratitude to the community-science volunteers of the Gravity Spy project, whose dedicated efforts played a pivotal role in bringing this project to fruition. The authors also thank ManLeong Chan and the anonymous referees for comments that improved the quality of this manuscript, and Shamal Lalvani, Yi Li and Jonathan Stromer-Galley for insightful conversations. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. Gravity Spy is supported by NSF grants IIS-2106882, 2106896, 2107334, and 2106865. CPLB acknowledges support from Science and Technology Facilities Council (STFC) grant ST/V005634/1. ZD acknowledges support from the CIERA Board of Visitors Research Professorship and NSF grant PHY-2207945. This document has been assigned LIGO document number LIGO-P2300458.
@@ Acknowledgments
The datasets from the Gravity Spy project, including machine learning and volunteer classification information, can be accessed through the Zenodo repositories [65, 40]. The corresponding codes are also accessible on GitHub through Gravity Spy.